OpenStudy (anonymous):

Simplify by using the distributive property. 10w-7w?, Please this is easy people, I just need confirmation

OpenStudy (anonymous):

3w?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

\[10w-7w = (10-7)w = 3w\]